OpenStudy (anonymous):

A muffin calls for 3/4 cups sugar for 3 dozen muffins. How much sugar will you use to make 8 dozen muffins?

OpenStudy (anonymous):

(3/4):3 x: 8 cross-multiply. (3/4)*8 = 3x 3x = 6 x=2

OpenStudy (anonymous):

thank you but I am still a little confused... :(

OpenStudy (anonymous):

hmm. well it's a ratio problem. So 3/4 cups of sugar gives 3 dozen. that means: (3/4) : 3 (<-- that's just the notation) so now you have a case where you have 8 dozen cookies. but you don't know how much sugar is needed. So let's make the amount of sugar needed to be x. x cups of sugar gives 8 dozen cookies. that means: x : 8 When you have this kind of ratio, you can cross multiply. So (3/4) * 8 = 3x. (3/4)*8 = 6. So 6 = 3x. x therefore equals 2 cups. hopefully that makes more sense. :)

OpenStudy (anonymous):

Thank you...I got the answer 2 but I did it a little different and just wasn't sure if it was just a accident...I multiplied 3/4 by 8/1 which gave me 24/4 which equals 6 and then divided by 3. which would equal 2.